### v3

# arXiv:2505.00784v3[cs.RO] 27 Feb 2026

arXiv:2505.00784v3[cs.RO] 27 Feb 2026
See pages - of Manuscript.pdf
Supplementary Information for
Agile legged locomotion in reconfigurable modular robots
Chen Yu1, David Matthews2, Jingxian Wang2, Jing Gu, Douglas Blackiston, Michael Rubenstein, Sam Kriegman
1,2Co-first authors.
2Order determined by coin toss.
To whom correspondence should be addressed.
Email: sam.kriegman@northwestern.edu
This PDF file includes:
Captions for Movies S1 to S3
Materials and Methods
Figures S1 to S9
Tables S1 to S10
Other supporting materials for this manuscript include the following:
Movies S1 to S3
S1 Supplemental Movies
Supplemental Movies S1–S3 can also be streamed online and are linked on our project page:
https://modularlegs.github.io
S1.1 Caption for Movie S1
Reconfigurable robots in the wild. Reconfigurable legged robots behave autonomously across sand, mud, grass, tree roots, plant litter, mulch, gravel, bricks, concrete, and combinations thereof.
S1.2 Caption for Movie S2
Acrobatic behaviors and resilience to damage. Robots behave, resist adversarial perturbations, and withstand body-altering damage. Control policies use internal sensing only; motion capture was not supplied to the policy, it is used only for the behavioral analyses in Figs. 2, 3 and 5 in the manuscript, as well as that of Figs. S1 and S2 below.
S1.3 Caption for Movie S3
Mechatronics of a modular leg. Mechatronic components are assembled into a modular leg.
S2 Single module hardware
The two links and central sphere of each module were 3D printed with PAHT-CF (Polyamide High-Temperature Carbon Fiber) filament on a Bambu Lab X1C 3D Printer. The PCB (Printed Circuit Board) cradle was printed with PLA (Polylactic acid) filament. Table S1 lists 3D printing parameters.
S2.1 PCB design
The PCB consists of an Espressif ESP32-S3-PICO-1-N8R2 microcontroller, Ceva BNO086 9-axis IMU, Qorvo DWM1000 Ultra-Wideband (UWB) module, Diodes Inc. AH49HNTR-G1 hall sensor, TDK ICS43434 microphone, and other supporting components (Fig. S6). The microphone and UWB module are not used in any of these experiments. The PCB design is open source and provided on our project webpage (https://modularlegs.github.io). The PCB is a four-layer design with a size of 95.6 $\times$ 60 mm, 1.6 mm thickness, and 1 oz copper traces.
S2.2 Communication
Each module has sufficient onboard processing capabilities for fully autonomous closed-loop control. However, for convenience we employ a remote computer to calculate module actions. Each module is connected to a remote computer through WiFi. The module publishes all the sensor data and a unique module identification number (bound to MAC address) to the computer over UDP at a frequency of 100 Hz. Sensor data used for observations include: IMU data and motor encoder data. Module reports additional sensing information (e.g. voltage, current, motor temperature, etc.) for logging purposes. The remote computer publishes motor position targets and PD (Proportional-Derivative) control parameters $K_{p}$ and $K_{d}$ to each module via UDP at 20Hz. Since the modules publish sensor data faster than the control frequency, the remote computer only utilizes the most recent data.
Each module independently sends observation and telemetry messages to the remote computer, and the remote computer uses the most recent message it receives for each module. If the remote computer fails to receive observation updates from one or more modules, then the action policy may take as input observations from modules at slightly different points in time. Note that use of UDP allows module messages to arrive out of order. Timing information was not provided to the modules, and the modules do not directly communicate with each other. While explicit synchronization is not necessary for the agile dynamic behaviors shown herein, the modules are fully equipped with hardware that can provide distributed and low-latency peer-to-peer communication (ESP-NOW, UWB).
S3 Single module controller
We model the control of the single module as a Partially Observable Markov Decision Process (POMDP), described by the tuple $(\mathcal{S},\mathcal{A},P,\mathcal{O},\mathcal{Z},\mathcal{R})$. The POMDP tuple consists of states $\bm{s}\in\mathcal{S}$, actions $\bm{a}\in\mathcal{A}$, state transition dynamics $P\left(\cdot\mid\bm{s},\bm{a}\right)$, observations $\bm{o}\in\mathcal{O}$, an observation model $\mathcal{Z}(\bm{o}\mid\bm{s},\bm{a})$, and a reward function $r=\mathcal{R}(\bm{s},\bm{a})$. We use $\bm{s}_{t}$, $\bm{o}_{t}$, $\bm{a}_{t}$, and $r_{t}$ to denote state, observation, action, and reward at timestep $t$, respectively.
S3.1 Observation space for single module controller
We add a cosine mapping to the joint position to deal with the general case where the joint can rotate continuously. All other details of the observation space are described as part of the Materials and Methods in the manuscript (under the heading, “Observation Space”).
S3.2 Reward function for single module rolling
[…]
${\omega}_{f}=\bm{\omega}\cdot\bm{\ell}^{\prime}$ (S1)
[…]
$r_{\text{roll}}=\min\left(1,\frac{\omega_{f}}{\omega^{\ast}}\right)$ (S2)
We penalize the action rate of all modules in a robot in simulation to encourage the smoothness of actions and help the sim-to-real transfer, as this term serves as a regulator in the policy optimization. The action rate reward at time step $t$ is defined as:
$r_{\text{action}}=\|\bm{a}_{t-1}-\bm{a}_{t}\|_{2}^{2}$ (S3)
The final reward is a linear combination of the forward reward and action rate penalty:
$r=\alpha_{\text{forward}}\,r_{\text{roll}}+\alpha_{\text{action}}\,r_{\text{action}}$ (S4)
with $\alpha_{\text{forward}}=1$ and $\alpha_{\text{action}}=-0.1$.
S3.3 Reward function for single module turning
The reward function for turning a single module is also defined by two terms: turning reward $r_{\text{turn}}$ and action rate penalty $r_{\text{action}}$. The turning speed $\omega_{\text{turn}}$ is defined by the dot product of the projected gravity and the angular velocity in the module’s body frame. The turning reward
$r_{\text{turn}}=\min\left(1,\frac{\omega_{\text{turn}}}{\omega_{\text{turn}}^{\ast}}\right)$ (S5)
is the fraction of the desired turning speed, capped at 1. $\omega_{\text{turn}}^{\ast}=3\,\text{rad/s}$. The final reward are calculated as:
$r=\alpha_{\text{turn}}\,r_{\text{turn}}+\alpha_{\text{action}}\,r_{\text{action}}$ (S6)
with $\alpha_{\text{turn}}=1$ and $\alpha_{\text{action}}=-0.1$.
S3.4 Domain randomization for single module training
Control policies were trained simulation using the Mujoco simulator  (?). Domain randomization was employed to ensure behaviors optimized in simulation transferred with sufficient fidelity to the physical system. Mass distribution, mass, module geometry, friction, and motor parameters were randomized during training. Mass distribution was randomized by sampling the motor position offset and motor mass. Domain randomization parameters are shown in Table S2, where $m$ is the mass of the agent (single module or multi-module robot). We uniformly sample these values at each episode of the training process. We also add Gaussian noise sampled from $\mathcal{N}(0,\,0.2)\,$ to the observations and noise from $\mathcal{N}(0,\,0.1)\,$ to the actions. One of the sim-to-real transfer challenges was the latency of the WiFi communication between the remote computer that runs the RL policy and the module. To model this latency in simulation, the simulated module randomly switches between executing the current action output by the policy or the previous action, with equal probability.
S3.5 Training single module controllers
We use CrossQ  (?) for training both single modules and multi-module robots. CrossQ adopts the same overall update rules as Soft Actor-Critic  (?), with separate updates for the critic, the actor, and the entropy temperature. The critic is updated by minimizing the temporal-difference error between the predicted Q-value and a bootstrap target; the actor is updated to favor actions that achieve high Q-values; and the temperature parameter is adjusted to keep the policy entropy close to a target value. See Bhatt et al. (2024;  (?)) for more details. A relatively small policy network (256, 256) was used for sample efficiency alongside a larger critic network (1024, 1024). Policies were trained $10^{6}$ time steps.
S3.6 Single module sim2real
When deployed to physical hardware, PD control parameters $K_{p}=12$ and $K_{d}=0.4$ were used. Sometimes, particularly on uneven or deformable surfaces, the position of the module can inhibit rolling. If the module detects that it is unable to initiate rolling, a joint angle is randomly sampled. Once the module is rolling, its kinetic energy keeps it moving forward. Note that stochastic actions were unnecessary and not utilized by the rolling policy in Fig. 2.
S3.7 Single module jumping
Due to the nature of the module design, it is trivial to perform a jumping behavior. Therefore, we didn’t train an RL policy for the jumping policy; Instead, we use a uniform stochastic policy with joint angles between -3 to 3 radians in the wild to perform the jumping.
S4 Multilegged configurations
S4.1 Docking
Module docks were designed for two key traits: flexibility and strength. For flexibility, we wish modules to connect with each other in as many diverse positions without incurring interference between docks. This motivated us to create docks that could be placed along the module’s central sphere, and on the side and tips of its links. Each dock has 3-fold rotational symmetry, further increasing the flexibility of connection. The number and location of docks was optimized to ensure machine screw pathways from one dock do not interfere with another. As a result, docking three modules can not be mutually orthogonal and connect at a single point. One of the three links would need to slide down along another to dock to the nearest orthogonal spot. Indeed this is precisely the three module design discovered by BO (Fig. 1C-E). Overall, there are 18 docks distributed along a module’s surface, with few combinations that lead to interference.
Since we desire robots composed of these modules to perform powerful, athletic behaviors, inter-module connections through docks must be strong enough to endure stress in all directions. We thus designed the docks to be honeycomb-shaped and non-flat so that the external force and torque can be largely combated by the normal force on the docking surface and the combined torque of the normal forces. This design helps to relieve stress on connecting bolts and bolts’ contact points on the robot, reducing failures and increasing rigidity of the docks.
With the design choices described above, our modules can dock together to form a wide variety of morphologies. Consider two modules, $i$ and $j$, which are connected together via a dock to form a two-module design. We first assume that these two modules are not identical. Because any two docks can connect together, each module has 18 docking points, and docks have 3-fold rotational symmetry, we have $18\times 18\times 3$ possible configurations. However, due to interference, two parallel modules cannot dock through docks on the sides of links while remaining parallel. Each module has 12 such docking points, and so interference eliminates $12\times 12$ configurations.
In reality, the modules are identical, so a design formed by module $i$ and $j$ will be the same if module $i$ takes module $j$’s place and vice versa. As such, we need to divide the numbers we have obtained by 2 to count the number of unique configurations. However, if both modules are connected through the same dock, swapping $i$ and $j$ will result in an identical configuration even if the modules are not identical, and these cases should not be divide by 2. Each module has 18 docking points, and docks have 3-fold rotational symmetry, thus we have $18\times 3$ configurations where both modules are connected through the same dock. However, among these possibilities, 12 cases will lead to interference as explained above and should be left out. So, in total we have $18\times 3-12$ configurations where both modules are connected through the same dock. After accounting for these cases, in total there are $(18\times 18\times 3-12\times 12+18\times 3-12)/2=435$ unique two module design.
As the number of modules increases, interference between modules and the symmetry of the design becomes much more challenging to evaluate, so we will provide an estimation of the number of unique configurations instead. Let’s assume that there are $M$ unique configurations for an $N$ module design. When adding another module to the $N$ module design, we have $16N+2$ free docks on an $N$ module design and 18 docks on the added module, giving us a total of $(16N+2)\times 18\times 3\times M\approx 864NM$ possible configurations. Considering that swapping the added module with any module in the $N$ module design will result in an identical design, we approximately have $864NM/(N+1)$ unique designs. Using mathematical induction, we get a good estimate for the number of unique $N$ module designs to be $864^{(N-1)}/N$. This number would grow to approximately $10^{11}$ when $N=5$.
S4.2 Morphological optimization
To find good designs with high locomotive ability, we encoded the combinatorial configuration space into a continuous 8D latent space using a variational autoencoder (VAE; (?)). A VAE was used because it transforms the search problem into a continuous one, which simplifies optimization  (?), and because we have considerable experience using this particular encoding. We searched for good configurations within this compressed latent design space using Bayesian Optimization (BO). We used multilayer perceptrons (MLP) with rectified linear unit (ReLU) activation functions for both the VAE encoder and decoder networks. The size of each hidden layer of the VAE encoder was (512, 128, 64, 64, 16) and the size of each hidden layer of the VAE decoder was (64, 64, 128, 512).
S4.3 Representing different designs (VAE input)
To input the topologies of modules into a VAE, we represent an assembled robot as a tree, where each node is an attached module. Each configuration tree can be uniquely represented as a sequence of integers with each connection pair represented by four items: the module ID of the parent module, the position ID of the dock on the parent module, the dock ID of the child module, and the orientation ID of these two docks. In this way, an assembled design of $N$ modules can be represented by $4\times(N-1)$ integers. To encode designs with $M$ modules when $M<N$, the final $4\times(N-M)$ integers are set to a reserved value which denotes “not present”. We employed a one-hot encoding on each integer in the sequence, and then fed the concatenated result to the VAE.
The one-hot encoding represents the integer IDs as binary vectors, where only one element is set to 1 and all others are set to 0. Sixteen one-hot vectors fully specify an arbitrary configuration tree with up to five modules. Each of the one-hot vectors has 19 elements. This length ensured vectors were long enough to capture the valid range of IDs. The 16 one-hot vectors were concatenated into a single binary vector (with $16\times 19=304$ elements) and fed as input to the VAE. During reconstruction, if the decoded one-hot representation contains invalid IDs, they were replaced with the nearest valid alternative. This prevents, for example, attaching a second module to a dock already occupied by a sibling on the parent. If multiple empty tokens appear, or if an empty value occurs at an invalid position, the design was interpreted as the complete tree up to the first valid empty value.
S4.4 VAE training
We sampled $5\times 10^{5}$ configuration trees with at least two and no more than five modules by uniformly randomly selecting the number of modules between two to five (inclusive). For each module connection, we randomly pick the parent node and the connection link between the parent node and the child node. During the sampling, configurations that have self-collision when all joint positions are zero, were discarded.
S4.5 Bayesian Optimization
Because training time can vary greatly when evaluating configurations with differing numbers of modules, we used asynchronous Bayesian Optimization (BO), allowing multiple workers to comb the latent space for good configurations, in parallel asynchronously. We use the method in  (?) to penalize the acquisition function using information about configurations that are still under evaluation. BO was used to identify configurations that were able to locomote well. The performance of a design for the purpose of BO was taken to be the average accumulated reward across the final 10% of episodes. Candidate designs identified through this method underwent training on additional behaviors before being selected for assembly.
S4.6 Initial pose optimization parameters
During BO, the initial pose of a selected configuration was optimized by sampling random poses and scoring them by their average speed under open loop control. Falling over during this preliminary assessment results in a penalty of -100 applied to the pose score.
S4.7 Promoting symmetrical poses
Inspired by legged animals, which lift themselves above the ground with pairs of legs, some BO trials augmented the pose heuristic function to promote taller, bilaterally symmetrical configurations. To do so, the pose of the root module was constrained to an initial joint angle of 0. Selection pressure for pairs of legs—modules docked in plane along the root module—was then applied by discarding configurations with more than one unpaired non-root module. The initial pose of leg pairs was constrained such that they shared the same absolute joint angle value. Also, in this initial pose, no joint spheres were permitted to contact the floor. These constraints led to the discovery of the three-module design featured the main manuscript (Fig. 4N). The other two designs discovered by BO (Fig. 4P and Fig. 4R) did not undergo this additional selection pressure.
S4.8 Advantages and limitations of the encoding
The VAE transforms the discrete space of configuration trees into a smaller, continuous latent space. This enables the use of any continuous optimization algorithm. One downside of the VAE approach, however, is that there is no guarantee that a particular design can be found or that it even exists in the latent space. For example, we were unable to identify latent coordinates corresponding to the manually designed quadruped. Passing the five-module quadrupedal configuration through the optimized encoder into latent coordinates, and then propagating those coordinates back through the optimized decoder does not fully reconstruct the original quadruped. The reconstruction contained five modules, but its forelimbs were shifted posteriorly along the spine, and its spine was elongated by an additional module, leaving a single hind limb.
Here, an important question is whether or not the compression generates latent representations that are more easily searchable than the uncompressed “raw” representation of the configuration tree. In other words, does the latent space facilitate the discovery of better designs with higher performance? To investigate this, BO was applied directly to the one-hot representation of configuration trees (see Sect. S4.3). Thus, BO directly outputs 16 probability vectors that, when converted into 16 one-hot channels, fully specify an arbitrary configuration tree with up to five modules. When using a VAE, these 16 one-hot channels were fed as input to the VAE. Without the VAE, these 16 one-hot channels were directly specified by BO. The results of this experiment are reported in Fig. S9. They suggest that BO in the latent space yields designs with higher performance than BO in the raw configuration space.
S5 Multilegged robot controllers
Similar to the single module training, we also model the control problem as a POMDP.
S5.1 Observation space for walking
The observation space of the assembled modules includes the projected gravity of the torso module $\bm{g}_{p}$, the angular velocity of the torso module $\bm{\omega}$, the joint angle $\bm{\theta}$ and joint angular velocity $\dot{\bm{\theta}}$ of every module, and the prior actions $\bm{a}_{\text{prior}}$. The observation consists of these signals from the prior three consecutive timesteps. We add a cosine mapping to the joint position to handle the general case where the joint can rotate continuously.
S5.2 Reward function for walking
Inspired by  (?), the reward of the walking task is a linear combination of six terms: forward velocity tracking reward $r_{\text{forward}}$, angular velocity tracking reward $r_{\text{walk-turn}}$, action rate penalty $r_{\text{action}}$, joint falling penalty $r_{\text{fall}}$, motor velocity penalty $r_{\text{motor-vel}}$, and joint acceleration penalty $r_{\text{motor-acc}}$. The forward velocity tracking reward is calculated by
$r_{\text{forward}}=\exp\left(-\frac{(v^{\ast}-v_{\text{forward}})^{2}}{\sigma_{\text{forward}}}\right),$ (S7)
where $v_{\text{forward}}=\mathbf{v}\cdot\hat{\bm{x}}$ is the projection of velocity in direction $\hat{\bm{x}}$, $\mathbf{v}$ is the robot’s velocity vector, $\hat{\bm{x}}$ is the unit local forward direction vector given by the initial pose from the pose optimizer, $v^{\ast}=0.6\,\text{m/s}$ is the desired forward speed, and $\sigma_{\text{forward}}=0.15$ is a hyperparameter. The angular velocity reward is calculated by
$r_{\text{walk-turn}}=\exp\left(-\frac{(\omega^{\ast}_{z}-\omega_{z})^{2}}{\sigma_{\text{walk-turn}}}\right),$ (S8)
where ${\omega}^{\ast}_{z}$ is the desired angular velocity around the global z axis, which is set to 0 here, and the ${\omega}_{z}=\bm{\omega}\cdot\bm{g}_{p}$ is the z component of angular velocity in the global frame; $\bm{\omega}$ is the robot’s angular velocity and $\bm{g}_{p}$ is the projected gravity. The hyperparameter $\sigma_{\text{walk-turn}}$ is set to 0.15. The action rate penalty is calculated in the same way as training for the single module as in Eqn. S3. The joint falling penalty $r_{\text{fall}}$ is the number of joints (the sphere part of the module) touching the floor for each time step. The joint velocity penalty is calculated by
$r_{\text{motor-vel}}=\sum_{i}\big(\big|\dot{\theta}_{i}\big|-\dot{\theta}_{\text{limit}}\big),$ (S9)
which will be clipped from 0 to $1\times 10^{5}$; $\dot{{\theta}}_{i}$ is the joint velocity of joint $i$ and $\dot{\theta}_{\text{limit}}$ is the joint velocity limit whose value is 10 radians/s in our experiment. The joint acceleration penalty is calculated by
$r_{\text{motor-acc}}=\left\lVert\frac{\dot{\bm{\theta}}_{\text{prior}}-\dot{\bm{\theta}}}{\text{dt}}\right\rVert_{2}^{2},$ (S10)
where $\dot{\bm{\theta}}$ is the velocity of all joints and $\dot{\bm{\theta}}_{\text{prior}}$ is the joint velocity at the prior time step. As the control frequency is 20 Hz, $\text{dt}=0.05s$ here. The reward weights are shown in Table S3.
S5.3 Action space for walking
Similar to the single module training, the action output by the policy, $\bm{a}_{t}$, is offset by predefined values for each joint. Before feeding the offset action to the PD controller, we filter the action with a Butterworth filter (high cutoff frequency is 3; low cutoff frequency is 0; filter order is 2) to promote the smoothness of the actions  (?). The action is then clipped to $\pm 1.2$ radians and sent to the PD controller. The torque output by the PD controller is again clipped by the TN (torque-speed) curve of the real motor.
S5.4 Termination of walking policy
Although we have a fixed maximum episode length, we terminate each episode earlier if and when the robot falls over. Inspired by  (?), falling is determined by the deviation of the body’s upward position $d$, which is defined by:
$d=-\bm{g}_{p}\cdot\hat{\bm{z}}$, (S11)
where $\bm{g}_{p}$ is the projected gravity and $\hat{\bm{z}}$ is the local unit upward vector given by the initial pose from the pose optimizer. The falling is determined when $d<\epsilon$, where $\epsilon$ is the threshold for falling detection, whose value is 0.1.
S5.5 Self-righting when inverted
After the design is optimized by BO, we retrain the optimized configuration with the optimized pose for other tasks like recovering an upright pose after being inverted. When falling is detected, the walking policy automatically switches the self-righting policy, which is activated for a fixed amount of time $T_{\text{activate}}$. The activation time $T_{\text{activate}}$ varies across different configurations. For the three-module design, $T_{\text{activate}}=1.5s$; for the four-module design, $T_{\text{activate}}=5s$, for the five-module design, $T_{\text{activate}}=3s$, for the example five-module quadrupedal design, $T_{\text{activate}}=3s$.
S5.6 Reward function for self-righting
The reward function for the self-righting policy consists of two terms: a pose reward $r_{\text{pose}}$ and an action rate penalty $r_{\text{action}}$. The pose reward $r_{\text{pose}}$ is defined as:
$r_{\text{pose}}=r_{\text{upward}}\,r_{\text{joint}}\\$ (S12)
where $r_{\text{upward}}$ measures how well the design stays in an upright position, defined as $r_{\text{upward}}=d$, where $d$ is the deviation of the body’s upward position defined in Eqn. S11. The pose reward $r_{\text{joint}}$ measures how close the current joint positions $\bm{\theta}$ are to the initial joint positions $\bm{\theta}_{0}$:
$r_{\text{joint}}=\exp\left(-\frac{\|\bm{\theta}_{0}-\bm{\theta}\|^{2}_{2}}{\sigma_{\text{joint}}}\right).$ (S13)
The action rate penalty is calculated as in Eqn. S3. The weight for the pose reward term is 1 and the weight for the action rate is 0.02-0.02. The hyperparameter $\sigma_{\text{joint}}$ here is set to 10.
S5.7 Action space for self-righting
To achieve dynamic self-righting behavior, we increase the action clipping from $\pm 1.2$ to $\pm 3.14$ radians.
S5.8 Observation space for jump turn
The observation space is the observation space for walking plus an additional dimension for a jumping command dimension, similar to  (?). The jumping command is set to 0 except when the robot is commanded to jump. In simulation, the jumping command is randomly sampled as 0 or 1 every 100 timesteps during training and is automatically reset to 0 when the robot reaches the desired height. After deploying the policy in the real world, the jump command is nominally 0 and can be set to 1 by a user. Since we assume only proprioceptive sensing on the physical hardware, rather than using a state estimator as in  (?) to measure the height of the robot for resetting the command, we set a fixed activation period $T_{\text{activate}}$ for the jumping command. $T_{\text{activate}}$ is set to 0.75 seconds in our experiment.
S5.9 Reward function for jump turn
The rewards of the jumping task consist of five terms: upward reward $r_{\text{upward}}$, pose reward $r_{\text{pose}}$, height track reward $r_{\text{height}}$, jump bonus $r_{\text{jump}}$, and turn reward $r_{\text{jump-turn}}$. The upward reward $r_{\text{upward}}$ is to prevent the robot from falling over. It is defined in the same way as in Eqn. S12. The pose reward $r_{\text{pose}}$ is to encourage the robot to retain the joint angles from the initial pose when the jumping command is not triggered. The pose reward is set to 0 when the jump signal is triggered (the jumping command in the observation is set to 1); when the jump signal is not triggered, the pose reward is defined as in Eqn. S12. The height track reward is to encourage the robot to jump to a desired height when the jumping command is triggered. When the jump signal is not triggered, $r_{\text{height}}=0$, otherwise
$r_{\text{height}}=\min(h,h^{\ast})$ (S14)
where $h$ is the height of the torso module and $h^{\ast}$ is the desired jumping height. The values of $h^{\ast}$ for each design can be found in Table S4. To further encourage the robot to jump, a bonus $r_{\text{jump}}$ is used, which is set to 1 when the jump signal is triggered and the robot is not touching the ground; otherwise $r_{\text{jump}}=0$. The turn reward $r_{\text{jump-turn}}$ is to encourage the robot to turn when jumping. When the jump signal is triggered,
$r_{\text{jump-turn}}=\min({\omega}_{z},\omega^{\ast}_{\text{jump}}),$ (S15)
where $\omega^{\ast}_{\text{jump}}=2\,\text{rad/s}$, otherwise $r_{\text{jump-turn}}$ is defined in the same way as $r_{\text{walk-turn}}$ in Eqn. S8. The reward weights are shown in Table S5.
S5.10 Action space for jump turn
To achieve dynamic jumping actions, we increase the action clipping to $\pm 2.5$ radians from $\pm 1.2$ radians for walking.
S5.11 Curriculum for policy training
There are two types of curriculum learning in the training process. Firstly, during the training of the agent in the evaluation period of BO, we adjust the weights of the reward function to guide a candidate design first to learn to walk and then to walk in a straight line. Secondly, after the design optimization process and before transferring the policy to the real world, we fine-tune the policy with a wider range of domain randomization to close the sim-to-real gap. For the first type of curriculum learning, we set the weight of the angular velocity reward term as 0.2 for the first $2\times 10^{5}$ steps and change it to 0.4 for the rest of the training (another $2\times 10^{5}$ steps). The beginning lower weight of the angular velocity reward term is to prevent the design from getting stuck at a local maximum. One such local optimum is maximizing the action smoothing and making the angular velocity close to zero just by standing still. Increasing the weight is to encourage the design to walk in a straight line. For the second type of curriculum learning, we train the design with a relatively low range of domain randomization to also prevent the design from getting stuck at a local maximum of not moving. The range of domain randomization in each phase is shown in Table S6. The parameters that are not shown in Table S6 have the same values as in Table S2.
S6 Amputation-agnostic control
We model the amputation-agnostic control problem as a POMDP. We desire a single policy that can control a robot after arbitrary amputations. We define the state space of the robot $\mathcal{S}_{\text{meta}}$ as the collection of the state space of each module $\mathcal{S}_{\text{module}}$ in this system:
$\mathcal{S}_{\text{meta}}=\mathcal{S}_{\text{module}}^{0}\cup\mathcal{S}_{\text{module}}^{1}\cdots\cup\mathcal{S}_{\text{module}}^{N},$ (S16)
where $\mathcal{S}_{\text{module}}^{i}$ is the state space of the $i$-th module and $N$ is the maximum number of modules in the system. The state space of each module includes projected gravity, angular velocity, cosine of the joint position, and the joint velocity. The action of this POMDP is the joint position of each module, and the reward function varies across the different reduced configurations (example amputations) used in training the amputation-agnostic policy (a more detailed explanation is provided below). We further reduce this POMDP to a sequence modeling problem  (?). Similar to  (?), we use a Transformer model  (?) to distill a general controller from several amputation-specific teacher controllers. However, rather than using a Transformer encoder as a student policy to directly map states to actions, we use a causally masked transformer decoder to fit a sequence of states and actions. By conditioning on the current state $\mathcal{S}_{\text{meta}}$, the transformer should predict the optimal action.
S6.1 Trajectory representation
Specifically, we regard the state of each module as a token and limit the maximum number of modules in the system to $N=5$. The trajectory representation is then:
$\tau=(\bm{s}^{0}_{0},\bm{s}^{1}_{0},\bm{s}^{2}_{0},\cdots,\bm{s}^{N}_{0},\bm{a}_{0},\bm{s}^{0}_{1},\cdots,\bm{s}^{N}_{1},\bm{a}_{1},\cdots,\bm{s}^{N}_{T},\bm{a}_{T}),$ (S17)
where $\bm{s}^{i}_{j}$ is the state of the $i$-th module at the $j$-th timestep and $\bm{a}_{j}$ is the single-token action for all $N$ modules at timestep $j$.
S6.2 Model architecture
For a time step $t$, we feed the most recent $K$ timesteps ($K=60$ in the experiment) of module states and actions into the Transformer, for a total of $K\times(N+1)$ tokens. Each of the module states and actions will be fed into its dedicated linear layer with layer normalization. We then add a positional embedding to the embedding of each token. Similar to Decision Transformer  (?), rather than using a standard positional embedding, we use the embedding of each timestep as the positional embedding (each of the $N+1$ tokens thus have the same positional embedding). The embeddings are then fed to a Generative Pre-trained Transformer (GPT) model  (?), which predicts the next action tokens through autoregressive modeling. The hyperparameters of the GPT model are shown in Table S9.
S6.3 Training data
A dataset of offline trajectories is given to train this model. In our experiment, we use the training dataset in Table S7. To generate this dataset, scenario-specific expert policies were trained from scratch to control the remnant structures. For training remnant configurations with more than one module remaining, we use the same observation space, action space, and reward function as the walking policy described above. For training a single module, we use the same observation, reward, and action as the single module rolling policy described above. However, when training against amputation examples, we add more randomization to the environment compared to the standard training methods described above. In the undamaged scenario, we randomize the length of all the limbs: Each link touching the floor is either with its nominal length or a half of its nominal length. In the four-remaining-modules scenario (one limb is removed), the farthest link of the limb diagonally across from the amputated limb is randomly assigned either its nominal length or 1.5 times its nominal length. For example, if the configuration is that the left forelimb is amputated, the right hind limb will have two possible lengths: the nominal length or an extended length. This encourages the robot to recover to an upright position. Training settings for single limb amputations of the quadruped are shown in Table S8. In the three-remaining-modules scenario (two limbs removed), we randomize the length of the forelimbs and the torso: Each link touching the ground is either with its nominal length or a half of its nominal length. When one module remains post-amputation, the left or right link is half of its nominal length. These remnant single modules were also trained on uneven terrain featuring obstacles of random heights. The heights of the obstacles are sampled from 0 to 0.03 m.
Note that the observation space of the POMDPs for generating the training dataset is different from the state space $\mathcal{S}_{\text{meta}}$. When collecting the rollouts after these policies are trained, the state of the robot $\bm{s}_{\text{meta}}\in\mathcal{S}_{\text{meta}}$ is collected rather than the observation of the specific amputation example. In addition, although the different amputation-specific policies have different zero positions for joints, we normalize them to have the same zero position when recording the rollout trajectories after training them. The training of each robot takes $1\times 10^{6}$ timesteps. To avoid the robot getting stuck at a local optimum and stopping motion, we set the weight of the forward term in the reward function as 1 at the beginning of training and tuned it down to 0.8 at the time step of $2\times 10^{5}$.
For a system of $N$ modules, we can merge any set of trajectories where the total number of modules is $N$. In our experiment, we take: one trajectory from one-module and four-module policy; one trajectory from a five-module policy; two trajectories from one-module policy and one trajectory from three-module policy; and five trajectories from one-module policy, to generate five-module training data trajectories.
S6.4 Model training
During training the amputation agnostic policy, batches of trajectories ($K$ timesteps of states/action pairs) from all amputation scenarios were sampled to train a GPT model. The prediction head that corresponds to the last module state token is trained to predict the action at that timestep with mean-squared error.
S6.5 Topologically in-distribution testing
At test time, we can specify the states of all modules as the conditioning information to generate the predicted best actions.
In Fig. 5, four scenarios are reported: no damage (all five modules intact), one module removed (four remaining), two modules removed (three remaining), and four modules removed (one remaining). Let an amputated topology be defined as the specific set of amputated modules. In these four scenarios, the four modules remaining case encapsulates four separate amputated topologies (front-right removed, front-left removed, back-right removed, or back-left removed). The other three scenarios each correspond to a single amputated topology (three modules remaining: loss of hindlimbs; and a single module remaining: loss of all limbs).
Recall that during training an amputated module is fully removed; however, during testing the precise location of amputation is uniformly sampled between the conjunction of the torso and each limb and the sphere of that limb. To quantify the performance of the amputation-agnostic policy compared to the original controller across previously unseen amputation locations, 10 sets of cutpoints for each amputation topology were sampled. Each set of remnant configurations was simulated for five seconds, five independent times under both the amputation-agnostic and original control polices. In Fig. 5H-K, the mean and standard deviation of displacement of each amputation scenario is reported across all constituent trials (H: no damage, I: four modules remaining, J: three modules remaining: K: one module remaining). The trajectories of the first of the five simulation repetitions for each of the sampled post-amputation robots under the amputation-agnostic policy (blue to green traces) and the original quadruped controller (pink to cyan traces) are shown in Fig. S3 (two rows per amputated topology). In most cases, the amputation-agnostic policy outperforms the original quadruped walking policy. The parameters of the testing simulation are set to match the real-world experiment, which are shown in Table S10.
S6.6 Topological out-of-distribution testing
To evaluate the amputation-agnostic policy against out-of-distribution amputations, we consider two new scenarios: two modules removed (three remaining), and three modules removed (two remaining). Unlike the in-distribution amputation tests, the out-of-distribution two modules removed class includes diagonal removal of one hindlimb and one forelimb (front-right and back-left, or front-left and back-right). The three modules removed case includes four separate amputated topologies (all limbs removed except any one of: front-right, front-left, back-right, back-left). For each amputated topology, trials are collected identically to Sect. S6.5, and the first of five simulation replications is plotted under the amputation agnostic policy (blue to green traces) and the original quadruped controller (pink to cyan traces) in Fig. S4 (two rows per amputated topology). Against these out-of-distribution amputations, the amputation-agnostic policy almost always outperforms the baseline controller.
S6.7 Dead modules (out of distribution)
Finally, we test the amputation-agnostic policy on another out-of-distribution scenario in which some modules are present but have died (disabled but still attached to the body). In this scenario, the output torques of one or two modules are set to zero to simulate module failure. Results are shown in Fig. S5, with dead modules colored red using the same blue-to-green and pink-to-cyan color maps as in Figs. S3-S4. Despite never experiencing dead modules during training, the amputation-agnostic policy generalized reasonably well to these scenarios and almost always outperformed the original quadruped controller.
S7 Supplemental Figures S1-S9
Figure S1: Jump Turn. The three (A-F), four (G-L) and five module (M-R) designs discovered by BO execute their jump-turn policies. Yaw, height, and joint positions are plotted during behavior.
Figure S2: Self righting. The three (A-F), four (G-L) and five module (M-R) designs discovered by BO were unexpectedly inverted by a wooden plank (red dot in B, H, and N). The designs rapidly contort and twist their body around to recover their upright poses and gaits (A-F, G-L, and M-R). Degrees of deviation from upright and joint positions are plotted during behavior.
Figure S3: The amputation-agnostic policy in different scenarios. Ten previously-unseen amputations, randomly sampled from each training scenario (one limb removed, two hindlimbs removed, all four limbs removed). For testing, amputated modules are partially removed which is distinct from from training where they were fully removed. Across nearly all tested scenarios, the amputation-agnostic policy (blue to green trajectories) substantially outperforms the original policy (cyan to pink trajectories). Denoted below each scenario is the locomotion speed (in m/s) generated by the original (dark purple; bottom left) and amputation agnostic (green; bottom right) policies.
Figure S4: The amputation-agnostic policy against out-of-distribution scenarios. The robustness of the amputation-agnostic policy is evaluated with respect to topologically distinct amputation classes composed of previously unseen sets of amputated modules. For each of six sets of module amputations (two diagonal limbs, and three limbs removed), ten amputation locations were randomly sampled (two rows of robots). Across nearly all damage scenarios, the amputation-agnostic policy (blue to green trajectories) substantially outperformed the original policy (cyan to pink trajectories). Denoted below each scenario is the locomotion speed (in m/s) generated by the original (dark purple; bottom left) and amputation agnostic (green; bottom right) policies.
Figure S5: The amputation-agnostic policy against out-of-distribution module failure. To mimic a realistic case of motor failure, another class of out-of-distribution damage scenario was considered. The torque output of one or two joints (highlighted in red) on a five-module design is zeroed resulting in a passive joint. The amputation-agnostic policy generally generalized reasonably well to these previously unexperienced failure mode, outperforming the original policy in most cases.
Figure S6: Front (A) and back (B) of PCB. The PCB consists of an Espressif ESP32-S3-PICO-1-N8R2 microcontroller, Ceva BNO086 9-axis IMU, Qorvo DWM1000 Ultra-Wideband (UWB) module, Diodes Inc. AH49HNTR-G1 hall sensor, TDK ICS43434 microphone, and other supporting components.
Figure S7: Overview of the design pipeline. Five hundred thousand example designs were randomly generated. Each design is a different configuration tree of at least three and no more than five modules, and is collision-free in its natural pose. A Variational Autoencoder (VAE) was trained to compress this dataset into a latent multivariate normal distribution, characterized by a mean vector ($\mu$) and standard deviations ($\sigma$). Training of the VAE minimizes two objectives: reconstruction loss, which measures how well decoded designs match the originals, and Kullback-Leibler (KL) divergence loss, which regularizes the latent distribution toward a standard normal prior. Bayesian optimization (BO) was conducted within the resulting latent space. BO builds a probabilistic surrogate model (Gaussian process) of the performance landscape, and uses this to propose which latent vectors to evaluate next, i.e. the latent coordinates that are predicted to be within promising unexplored regions of the latent space. Each latent vector is decoded into a configuration tree, instantiated in a physics simulation, and trained to locomote using deep reinforcement learning (RL). The performance of the design was taken to be the average accumulated reward across the final 10% of training episodes. This performance score is then fed back to BO to update the Gaussian process. In RL, an Actor network generates actions from states, while a Critic network estimates Q-values (expected future rewards) to update the Actor network. Optimized Actor policies were deployed on real robots for zero-shot control.
Figure S8: Overview of amputation-agnostic policy training and testing. In each training scenario, different combinations of limbs were amputated from the quadruped. For each scenario, a scenario-specific “expert” policy was trained to control the reduced morphology using reinforcement learning (RL). The sensorimotor sequences generated by the expert policy were collected for each scenario and pooled across scenarios in order to train a single transformer-based controller, an amputation-agnostic policy that generalizes across amputations. At test time, the amputation-agnostic policy was evaluated on diverse, unseen damage scenarios: novel amputation topologies, novel cut positions, and unexpected motor failures.
Figure S9: The utility of the latent design space. In the main experiment, bayesian optimization (BO) operates within a learned latent compression of configuration space (Fig. S7). To determine if this latent space is more easily searchable than the raw (uncompressed) configuration space, an otherwise equivalent experiment was conducted in which BO operates directly on the configuration trees (their one-hot encoding) instead of within the latent space (see Sect. S4.8 for details). Twelve independent trials of BO were conducted, six in the latent space (blue) and six on the raw configuration trees (red). In each trial, the performance of the best design found so far (i.e. cumulative max performance) was recorded across optimization time. These data are summarized with shaded regions indicating the standard error of the mean. On average, BO in latent space discovered significantly better designs than BO in the raw configuration space ($p<0.01$; t test). The performance of the manually designed quadruped is also shown for reference (gray dashed line). In all 12 trials, BO discovered designs with higher performance than the quadruped.
S8 Supplemental Tables S1-S10
Table S1: 3D Printing Parameters
Parameter Links Battery-side Motor-side PCB cradle Socks
hemisphere hemisphere
Material PAHT-CF PAHT-CF PAHT-CF PLA 95A TPU
Layer height [mm] 0.18 0.12 0.18 0.12 0.3
Line width [mm] 0.68 0.42 0.62 0.42 0.62
Wall loops 2 (5 at the tip) 2 3 (5 at the center) 5 2
Infill density 10% 10% 10% 10% 10%
Infill pattern Gyroid Gyroid Gyroid Gyroid Gyroid
Table S2: Domain Randomization of a Single Module
Parameters Lowest value Highest value
Motor mass [kg] 0.45 0.6
Total mass 0.7mm 1.3mm
Friction coefficient 0.6 0.8
Kp 4 8
Kd 0.1 0.3
Leg length [m] 0.22 0.28
Motor position offset [m] 0.0025-0.0025 0.0025
Joint damping 0.02 0.06
Joint armature 0.01 0.02
Table S3: Walking Reward Terms
Term Symbol Weight Weight (symmetrical configurations)
Forward velocity $r_{\text{forward}}$ 0.8 0.6
Angular velocity $r_{\text{walk-turn}}$ 0.2 0.2
Action rate $r_{\text{action}}$ 0 0.1-0.1
Num joint fall $r_{\text{fall}}$ 0.01-0.01 0.02-0.02
Vel penalty $r_{\text{motor-vel}}$ 0.005-0.005 0.01-0.01
Acc penalty $r_{\text{motor-acc}}$ 0.000001-0.000001 0.000002-0.000002
Table S4: Desired Height of Jump Turn
Metamachine Desired Height $h^{\ast}$
Five-module quadrupedal design 0.6 m
Three-module BO design 0.5 m
Four-module BO design 0.8 m
Five-module BO design 0.7 m
Table S5: Jump Turn Reward Terms
Term Symbol Weight
Upward $r_{\text{upward}}$ 0.2
Pose $r_{\text{pose}}$ 1
Height track $r_{\text{height}}$ 1
Jump bonus $r_{\text{jump}}$ 100
Turn $r_{\text{jump-turn}}$ 1
Table S6: Metamachine Domain Randomization Curriculum
Parameter Range (Phase 1) Range (Phase 2)
Friction [0.8, 1.2] [0.4, 0.8]
Mass [0.9mm, 1.1mm] [0.8mm, 1.2mm]
Table S7: Amputation-agnostic Policy Training Data
Configuration Total timesteps
Single module remaining 99999405
Three modules remaining 39999802
Four modules (front-right removed) 29999005
Four modules (front-left removed) 29996704
Four modules (back-right removed) 59995909
Four modules (back-left removed) 59996913
Quadruped (all five modules intact) 39997005
Total 359984743
Table S8: Training Setup for Quadruped with One Limb Amputated
Amputation site front-right front-left back-right back-left
Initial torso joint position [radian] 0.6 0.6-0.6 1 1-1
Table S9: Amputation-agnostic Transformer Architecture
Parameter Value
Context length, $K$ 60
Number of blocks 12
Embedding dimensions 768
Number of heads 12
Table S10: Testing Simulation Environment Parameters
Parameter Value
Friction coefficient 0.8
Initial torso joint position 0
Initial limb joint position 1.5